OpenStudy (fatimaj13):

cos theta -cos theta sin^2 theta. I know the answer is cos ^3 theta but i'm not sure how i got the answer

OpenStudy (agent0smith):

Factor out a cos theta\[\large \cos \theta -\cos \theta \sin^2 \theta = \cos \theta (1 - \sin^2 \theta) = ?\]Hopefully you recognize an identity there...

jhonyy9 (jhonyy9):

we know that sin^2 theta +cos^2 theta = 1 so from this result 1-sin^2 theta = cos^2 theta using this and what above wrote @agent0smith hope you can solve it easy